# UAlign: Leveraging Uncertainty Estimations for Factuality Alignment on Large Language Models

### v1

@@ 2.1.1 Responses Sampling Strategy
As in Fig. 2, to explore the knowledge boundary of the LLM given a question, we sample multiple responses by repeating the generation procedure several times. In this phase, the preparation process can be represented in a tuple $\left(\mathcal{Q},\mathcal{P},\mathcal{A}\right)$. $\mathcal{Q}$ contains a batch of $N$ QA pairs ${\left\{(\boldsymbol{x}_{i},\boldsymbol{\hat{y}}_{i})\right\}}_{i=1}^{N}$ where $\boldsymbol{x}_{i}$ and $\boldsymbol{\hat{y}}_{i}$ denote the $i$-th question and ground-truth answer respectively. To mitigate context sensitivity, we utilize different few-shot prompts in $\mathcal{P}$ with temperature $T=0.2$ to make a trade-off between the accuracy and diversity to represent knowledge boundaries (Gekhman et al., 2024). The few-shot prompt set $\mathcal{P}$ consists of $K$ different 1-shot exemplars in this work which is enough for LLMs to generate answers in the correct format. We present the few-shot prompts for sampling on TriviaQA and SciQ datasets as exemplified in Appendix I.
In the $k$-th sampling process for the $i$-th question $\boldsymbol{x}_{i}$, we employ each few-shot exemplar $\boldsymbol{p}_{k}\in\mathcal{P}$ with the question $\boldsymbol{x}_{i}$ to the LLM to generate the $k$-th response ${\boldsymbol{y}_{i}}^{(k)}$. By taking $K$ times of the sampling process, we can obtain an answer set $\boldsymbol{Y}_{i}={\left\{{\boldsymbol{y}_{i}}^{(k)}\right\}}_{k=1}^{K}$ to $\boldsymbol{x}_{i}$. We set the labels $\boldsymbol{Z}_{i}={\left\{{z_{i}}^{(k)}\right\}}_{k=1}^{K}$ by comparing each generated answer ${\boldsymbol{y}_{i}}^{(k)}$ with the ground-truth $\boldsymbol{\hat{y}}_{i}$ to indicate the correctness (${z_{i}}^{(k)}\in\left\{0,1\right\}$, 1 for True and 0 for False). We collect and format the data in $(\boldsymbol{x}_{i},\boldsymbol{Y}_{i},\boldsymbol{Z}_{i},\boldsymbol{\hat{y}}_{i})$ in an extended dataset and calculate the uncertainty measures subsequently. Note that since fine-tuning LLMs on unknown knowledge will encourage hallucinations (Zhang et al., 2024a; Gekhman et al., 2024), we revise the ground-truth answer to the question with ${z_{i}}^{(k)}=0,\forall{z_{i}}^{(k)}\in\boldsymbol{Z}_{i}$ to “Sorry, I don’t known.” to teach LLMs to refuse the questions beyond their knowledge (Zhang et al., 2024a).
@@ 2.2.2 UAlign PPO: Policy Model Training
The UAlign PPO is to elicit the LLM’s factual expressions to a question with the uncertainty measures using obtained models. Inspired by the progress of reinforcement learning from human feedback (RLHF) technique Ouyang et al. (2022); Ziegler et al. (2019), we employ proximal policy optimization (PPO) (Schulman et al., 2017) for LLM optimization with the reward model $\theta$. As illustrated in Fig. 3 (b), the LLM to be optimized is used as the policy $\pi_{\theta}$. During this phase, we iteratively feed the question $\boldsymbol{x}$, and the predicted confidence $c$ and entropy $e$ to both the policy $\pi_{\theta}$ and the reference $\pi_{o}$, and the reward function $r$ will facilitate reliable expressions of $\boldsymbol{y}$ of the policy model $\pi_{\theta}$. The training objective is to maximize the following reward function $r$ as
@@ 3.1 Datasets
For testing, we evaluate the in-domain (ID) performance on the corresponding validation/test sets and generalization on an out-of-domain (OOD) test set LSQA (Xue et al., 2024) which contains multilingual language-specific QA pairs. More dataset details and statistics are presented in Appendix B.
@@ 3.2 Evaluation Metrics
To evaluate the reliability of LLMs, we employ two metrics: Precision (Prec.) and Truthfulness (Truth.). Precision is defined as the proportion of correctly answered questions among all the known questions, representing LLMs’ ability to accurately express their known factual knowledge. Truthfulness represents the proportion of the sum of correctly answered known questions and refused unknown questions among all questions, indicating the honesty level of the LLMs.
To ascertain the correctness of the LLM-generated answer $\boldsymbol{y}$ with the ground truth $\boldsymbol{\hat{y}}$, we employ a string-matching approach. Exact matching (EM) of $\boldsymbol{y}\equiv\boldsymbol{\hat{y}}$ always misjudges some correct answers with slight distinctions on such closed-book QA tasks. Therefore, we replace EM with a variant of $\boldsymbol{y}\in\boldsymbol{\hat{y}}\vee\boldsymbol{\hat{y}}\in\boldsymbol{y}$ to evaluate the accuracy. The specific illustrations of evaluation formulas and comparisons of several EM variants we tested with human evaluations are in Appendix C.
@@ 3.3 Baselines
We present several baselines in four categories below. To clearly delineate the differences between our proposed method and other baselines, we have illustrated all methods in Fig. 7 in Appendix D.
@@
We present two prompt-based baselines namely In-Context Learning (ICL), In-Context Learning with Refusal Examples (ICL-IDK), and In-Context Learning Chain-of-Thought (ICL-CoT) (Wei et al., 2022). The few-shot prompt templates are presented in Appendix E.
@@ 3.4 Implementation Details
Experiments are conducted on two LLMs: Llama-3-8B (Llama-3)  [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B (AI@Meta, 2024) and Mistral-7B (Mistral)  [^3] https://huggingface.co/mistralai/Mistral-7B-v0.1 (Jiang et al., 2023). When preparing the UAlign dataset, we sample 10 responses for each question on $K=10$ different 1-shot prompts. The sampling temperature $T$ is set to 0.2 to achieve a trade-off between the diversity and factuality of the answer set. During training, all the LLMs are trained using LoRA (Hu et al., 2022) with rank $r=16$. Both the uncertainty estimation models and the reward model utilize the vanilla LLM as their bases and are trained using LoRA with rank $r=4$. ADAM parameter update is used in a mini-batch mode. Uncertainty estimation models and the reward model are trained using SFT on the UAlign dataset. The UAlign PPO algorithm and all the RL-based baselines are implemented by trl  [^4] https://github.com/huggingface/trl. All training hyper-parameters are presented in Appendix F. When decoding, the temperature is also set to 0.2 to be consistent with the sampling setting. All the experiments are conducted on 4 $\times$ NVIDIA A100-40GB GPUs.
@@ 4.1 Main Experimental Results
Significant improvements are consistently achieved on diverse datasets using the proposed UAlign framework over other baseline methods on both Llama-3 and Mistral. We highlight the supreme Precision and Truthfulness performance using grey highlights among the all baselines of each column in Table 1. The core idea of our UAlign framework is the utilization of uncertainty estimation models. Compared with the most relevant baselines of RL-PPO and RLKF, both the reward model and policy model in UAlign generate predictions and responses conditioned on uncertainty estimations regarding the knowledge boundaries to questions, thereby yielding better reliability performance. It can be attributed that by explicitly appending uncertainty measures following the question, LLMs can assist LLMs in eliciting more accurate responses based on intrinsic knowledge boundary representations.
@@ 4.3 Reliability of Uncertainty Estimations
Evaluating the performance of confidence score and semantic entropy is essential to the UAlign method. We present the AUROC (Detailed in Appendix C) results of two estimations in comparison with three confidence/uncertainty estimation methods (one probability-based method (Prob.), two prompt-based methods including p(True) and verbalized (Verb.) as illustrated in Fig. 8) on TriviaQA on two LLMs. Results on other datasets are remained in Appendix H. Details of baseline estimation baselines are presented in Sec. 5, Appendix G, and Fig. 8.
@@ Results
Figure 6: Experiments of Precision (Prec.) and Truthfulness (Truth.) of various sampling number $K$ of 1, 4, 7, and 10 on TVQA on Llama-3 and Mistral.
@@ 4.4 Effects of Sampling Number
The sampling number $K$ is a crucial hyper-parameter in the UAlign method. Different values of $K$ can significantly affect the precision of the knowledge boundary measurements. To evaluate the effects, we compare performances using various $K$ of 1, 4, 7, and 10. Experimental results on TVQA are presented in Fig. 6 in Appendix H.
@@
The experiments indicate that when using small sampling numbers, increasing the $K$ leads to significant improvements in both precision and truthfulness. However, as $K$ continues to increase, the reliability improvement tends to plateau, exhibiting convergence. Since further increasing $K$ requires substantial computational costs, we discard conducting experiments with larger $K$.
@@ 5 Related Works
Previous works investigate the knowledge boundary to identify the known level of a knowledge piece of LLMs by quantifying the confidence or uncertainty estimations like output consistency (Cheng et al., 2024), prompting methods (Ren et al., 2023) or knowledge probing (Ji et al., 2024). Generally, knowledge boundary measures derive from uncertainty estimations.
@@
Due to the space limitation, detailed investigations of related works are shown in Appendix G.
@@ 6 Conclusion
Computational Resources:
The current method for constructing the UAlign dataset relies on multiple samplings, requiring substantial computational cost which linearly increases with the number of sampling instances $K$ and significantly constrains the scalability expansion of the dataset in this work. To accurately approximate the knowledge distributions, a higher number of samplings is typically more beneficial. Here, we may need to introduce some prior knowledge distributions to alleviate the computational resource requirements and reduce the costs.
Task Expansion:
@@ A.2 Terminology Use
Appendix B Dataset Details

### v2

@@ 2.1.1 Responses Sampling Strategy
As in Fig. 2, to explore the knowledge boundary of the LLM given a question, we sample multiple responses by repeating the generation procedure several times. In this phase, the preparation process can be represented in a tuple $\left(\mathcal{Q},\mathcal{P},\mathcal{A}\right)$. $\mathcal{Q}$ contains a batch of $N$ QA pairs ${\left\{(\boldsymbol{x}_{i},\boldsymbol{\hat{y}}_{i})\right\}}_{i=1}^{N}$ where $\boldsymbol{x}_{i}$ and $\boldsymbol{\hat{y}}_{i}$ denote the $i$-th question and ground-truth answer respectively. To mitigate context sensitivity, we utilize different few-shot prompts in $\mathcal{P}$ with temperature $T=0.2$ to make a trade-off between the accuracy and diversity to represent knowledge boundaries (Gekhman et al., 2024). The few-shot prompt set $\mathcal{P}$ consists of $K$ different 1-shot exemplars in this work which is enough for LLMs to generate answers in the correct format. We present the few-shot prompts for sampling on TriviaQA and SciQ datasets as exemplified in Appendix J.
In the $k$-th sampling process for the $i$-th question $\boldsymbol{x}_{i}$, we employ each few-shot exemplar $\boldsymbol{p}_{k}\in\mathcal{P}$ with the question $\boldsymbol{x}_{i}$ to the LLM to generate the $k$-th response ${\boldsymbol{y}_{i}}^{(k)}$. By taking $K$ times of the sampling process, we can obtain an answer set $\boldsymbol{Y}_{i}={\left\{{\boldsymbol{y}_{i}}^{(k)}\right\}}_{k=1}^{K}$ to $\boldsymbol{x}_{i}$. We set the labels $\boldsymbol{Z}_{i}={\left\{{z_{i}}^{(k)}\right\}}_{k=1}^{K}$ by comparing each generated answer ${\boldsymbol{y}_{i}}^{(k)}$ with the ground-truth $\boldsymbol{\hat{y}}_{i}$ to indicate the correctness (${z_{i}}^{(k)}\in\left\{0,1\right\}$, 1 for True and 0 for False). We collect and format the data in $(\boldsymbol{x}_{i},\boldsymbol{Y}_{i},\boldsymbol{Z}_{i},\boldsymbol{\hat{y}}_{i})$ in an extended dataset and calculate the uncertainty measures subsequently. Note that since fine-tuning LLMs on unknown knowledge will encourage hallucinations (Zhang et al., 2024a; Gekhman et al., 2024), we revise the ground-truth answer to the question with ${z_{i}}^{(k)}=0,\forall{z_{i}}^{(k)}\in\boldsymbol{Z}_{i}$ to “Sorry, I don’t know.” to teach LLMs to refuse the questions beyond their knowledge (Zhang et al., 2024a).
@@ 2.2.2 UAlign PPO: Policy Model Training
The UAlign PPO is to elicit the LLM’s factual expressions to a question with the uncertainty measures using obtained models. Inspired by the progress of reinforcement learning from human feedback (RLHF) technique Ouyang et al. (2022); Ziegler et al. (2019), we employ proximal policy optimization (PPO) (Schulman et al., 2017) for LLM optimization with the reward model $\theta$. As illustrated in Fig. 3 (b), the LLM to be optimized is used as the policy $\pi_{\theta}$. During this phase, we iteratively feed the question $\boldsymbol{x}$, and the predicted confidence $c$ and entropy $e$ to both the policy $\pi_{\theta}$ and the reference $\pi_{o}$, and the reward function $r$ will facilitate reliable expressions of $\boldsymbol{y}$ of the policy model $\pi_{\theta}$. Model update details are further specified in Appendix B.1. The training objective is to maximize the following reward function $r$ as
@@ 3.1 Datasets
For testing, we evaluate the in-domain (ID) performance on the corresponding validation/test sets and generalization on an out-of-domain (OOD) test set LSQA (Xue et al., 2024) which contains multilingual language-specific QA pairs. More dataset details and statistics are presented in Appendix C.
@@ 3.2 Evaluation Metrics
To evaluate LLMs’ reliability, we employ two metrics: Precision (Prec.) and Truthfulness (Truth.). Precision is defined as the proportion of correctly answered questions among all the known questions, representing LLMs’ ability to accurately express their known factual knowledge. Truthfulness represents the proportion of the sum of correctly answered known and refused unknown questions among all questions, indicating LLMs’ honesty level. Details can be referred to Appendix D.1.
To ascertain the correctness of the LLM-generated answer $\boldsymbol{y}$ with the ground truth $\boldsymbol{\hat{y}}$, we employ a string-matching approach. Exact matching (EM) of $\boldsymbol{y}\equiv\boldsymbol{\hat{y}}$ always misjudges some correct answers with slight distinctions on such closed-book QA tasks. Therefore, we replace EM with a variant of $\boldsymbol{y}\in\boldsymbol{\hat{y}}\vee\boldsymbol{\hat{y}}\in\boldsymbol{y}$ to evaluate the accuracy. The specific illustrations of evaluation formulas and comparisons of several EM variants we tested with human evaluations are in Appendix D.2.
@@ 3.3 Baselines
We present several baselines in four categories below. To clearly delineate the differences between our proposed method and other baselines, we have illustrated all methods in Fig. 7 in Appendix E.
@@
We present two prompt-based baselines namely In-Context Learning (ICL), In-Context Learning with Refusal Examples (ICL-IDK), and In-Context Learning Chain-of-Thought (ICL-CoT) (Wei et al., 2022). The few-shot prompt templates are presented in Appendix F.
@@ 3.4 Implementation Details
Experiments are conducted on two LLMs: Llama-3-8B (Llama-3)  [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B (AI@Meta, 2024) and Mistral-7B (Mistral)  [^3] https://huggingface.co/mistralai/Mistral-7B-v0.1 (Jiang et al., 2023). When preparing the UAlign dataset, we sample 10 responses for each question on $K=10$ different 1-shot prompts. The sampling temperature $T$ is set to 0.2 to achieve a trade-off between the diversity and factuality of the answer set. During training, all the LLMs are trained using LoRA (Hu et al., 2022) with rank $r=16$. Both the uncertainty estimation models and the reward model utilize the vanilla LLM as their bases and are trained using LoRA with rank $r=4$. ADAM parameter update is used in a mini-batch mode. Uncertainty estimation models and the reward model are trained using SFT on the UAlign dataset. The UAlign PPO algorithm and all the RL-based baselines are implemented by trl  [^4] https://github.com/huggingface/trl. All training hyper-parameters are presented in Appendix G. When decoding, the temperature is also set to 0.2 to be consistent with the sampling setting. All the experiments are conducted on 4 $\times$ NVIDIA A100-40GB GPUs.
@@ 4.1 Main Experimental Results
Significant improvements are consistently achieved on diverse datasets using the proposed UAlign framework over other baseline methods on both Llama-3 and Mistral. We highlight the supreme Precision and Truthfulness performance using grey highlights among the all baselines of each column in Table 1. The core idea of our UAlign framework is the utilization of uncertainty estimation models. Compared with the most relevant baselines of RL-PPO and RLKF, both the reward model and policy model in UAlign generate predictions and responses conditioned on uncertainty estimations regarding the knowledge boundaries to questions, thereby yielding better reliability performance. It can be attributed that by explicitly appending uncertainty measures following the question, LLMs can be assisted to elicit more accurate responses based on intrinsic knowledge boundary representations.
@@ 4.3 Reliability of Uncertainty Estimations
Evaluating the performance of confidence score and semantic entropy is essential to the UAlign method. We present the AUROC (Detailed in Appendix D) results of two estimations in comparison with three confidence/uncertainty estimation methods (one probability-based method (Prob.), two prompt-based methods including p(True) and verbalized (Verb.) as illustrated in Fig. 8) on TriviaQA on two LLMs. Results on other datasets are remained in Appendix I. Details of baseline estimation baselines are presented in Sec. 5, Appendix H, and Fig. 8.
@@ Results
Figure 6: Experiments of Prec., Truth. (left), and time costs (right) of various sampling number $K$ of 1, 4, 7, and 10 on TVQA on both Llama-3 and Mistral.
@@ 4.4 Effects of Sampling Number
The sampling number $K$ is a crucial hyper-parameter in the UAlign method. Different values of $K$ can significantly affect the precision of the knowledge boundary measurements. To evaluate the effects, we compare performances using various $K$ of 1, 4, 7, and 10. Experimental results on TVQA are presented in Fig. 6.
@@
Results in Fig. 6 indicate that when using small sampling numbers, increasing $K$ leads to significant improvements in both Prec. and Truth.. However, as $K$ increases, the reliability improvement tends to plateau, exhibiting convergence. Therefore, we opt $K=10$ as the optimal setting and don’t experiment using larger $K$.
We also report the sampling time costs to construct the training set in Fig. 6, and further specify cost analysis of UAlign dataset construction and inference time in Sec. 4.5, Appendix B.2 and B.3 respectively. We showcase that with various acceleration and quantization methods, time costs of UAlign can be significantly reduced when scaling to larger models or datasets, exhibiting both efficiency and efficacy.
@@
4.5 Computation Cost Analysis
We test the time costs to construct the UAlign dataset $\mathcal{D}$ in different sampling numbers. We present different sampling time costs on 10000 QA samples on Llama-3-8B on 4×40G A100 GPUs as presented in Fig. 6. Results demonstrate the efficiency of our proposed UAlign method even though scaling on larger models. The relatively low computation costs when sampling can be attributed that experiments are conducted on knowledge QA datasets. The answer spans are entity-level, and each answer only needs to generate a few tokens. Since the output form is relatively simple, sampling ten times is sufficient and cost-saving to accurately fit the knowledge boundaries as in Sec. 4.4.
Furthermore, Test Time Scaling Law (Snell et al., 2024) has attracted much attention recently, which proposed to consider allocating more computation resources in inference to generate high-quality responses. These LLMs’ self-generated data can be further used for LLM training to self-improve LLMs (Gulcehre et al., 2023). Many works validate that incorporating multiply-sampled data on LLMs in inference can benefit LLMs for further improvements, which is a new trend for LLM training. In this way, our proposed UAlign provides a novel insight into the test time scaling law to represent knowledge boundaries by calculating uncertainty estimations on the sampled responses, and further explicitly leverages the uncertainty estimations for factuality alignment, heralding a promising view of the test time scaling law. Although a few additional computation costs are required, our proposed UAlign is still efficient to be utilized practically with significant reliability improvements.
@@ 5 Related Works
Previous works investigate the knowledge boundary (Yin et al., 2024; Wang et al., 2025) to identify the known level of a knowledge piece of LLMs by quantifying uncertainty estimations like output consistency (Cheng et al., 2024), prompting methods (Ren et al., 2023) or knowledge probing (Ji et al., 2024). Generally, knowledge boundary measures derive from uncertainty estimations.
@@
Due to the space limitation, detailed investigations of related works are shown in Appendix H.
@@ 6 Conclusion
Task Expansion
@@
Computational Cost
The current method for constructing the UAlign dataset relies on multiple samplings, requiring additional computational cost that linearly increases with the number of sampling instances $K$ and a higher number of samplings is preferable to accurately approximate the knowledge boundaries. As we have adopted a range of acceleration and quantization methods to reduce the time cost during both constructing the dataset and inference as presented in Appendix B.2 and B.3, there remains potential for exploration to further alleviate computational resources requirements.
Ethical Statement
In this paper, three evaluators are employed to annotate the correctness of four EM variants on selected samples, which aims to select the optimal EM variant to evaluate the correctness of the generated answer and the ground-truth label as presented in Appendix D.2. All the evaluators are M.Phil. or Ph.D. students possessing sufficient expertise to carry out the evaluation. We meticulously adhered to legal and ethical standards throughout the human evaluation process, prioritizing privacy and obtaining informed consent. Evaluators were furnished with comprehensive details regarding the study’s objectives, data collection methodologies, and associated risks or benefits. They were afforded the opportunity to seek clarifications and voluntarily provide consent before their involvement. All the human evaluation results were exclusively utilized for research purposes.
@@ Acknowledgments
This work was partially supported by Hong Kong RGC GRF No. 14206324, CUHK direct grant No. 4055209, and CUHK Knowledge Transfer Project Fund No. KPF23GWP20.
@@ A.2 Terminology Use
Appendix B Method Specification and Supplement
B.1 Model Update during PPO
During the PPO process, only the policy model $\pi_{\theta}$ is optimized while the uncertainty models $\mu,\tau$ do not need to be updated because the reward model $\theta$ updates are offline. As discussed and demonstrated in Sec. 4.2 and Table 2, uncertainty models are directly associated with and benefit the reward model. In our UAlign PPO algorithm, by incorporating the two uncertainty estimations, the reward model $\theta$ can provide more precise reward scores, thereby guiding LLMs $\pi$ to generate more factual responses. Since the reward model is offline during PPO, the uncertainty models also do not require online updates.
In addition, due to the KL-divergence constraint, the knowledge distribution of policy LLMs may not diverge too much from the initial policy. Both uncertainty models and reward models are trained on data generated by sampling from the vanilla LLMs, and their combined effect is to elicit the LLMs’ capacity for factual expression, evolving towards improved reliability. During the PPO process, with the KL-divergence constraint in Equation 2.2.2, the knowledge distribution of policy LLMs may not shift too much from the initial policy. We demonstrate the accuracy-based confidence distribution of Llama-3 before and after UAlign training on TriviaQA validation sets as follows.
Conf. Range Before UAlign After UAlign
$[0,0.25)$ 2404 2116
$[0.25,0.5)$ 1786 1628
$[0.5,0.75)$ 1509 1747
$[0.75,1.0]$ 4261 4469
Table 4: Accuracy-based confidence distribution of Llama-3 before and after UAlign training on TriviaQA validation sets.
Since the knowledge distribution does not shift too much from the initial policy, we can still achieve good performance without updating the uncertainty model for simplicity. Compared to traditional RLHF that solely utilize reward models, our proposed UAlign introduces uncertainty models that leverage knowledge boundary representations to benefit reward model and finally enhance LLMs, leading to signification improvements in reliability and generalization of knowledge QA tasks.
B.2 Computation Cost of Constructing UAlign Dataset
As mentioned in Sec. 4.4, we test the time costs to construct the UAlign dataset $\mathcal{D}$ in different sampling numbers. We present different sampling time costs on 10000 QA samples on Llama-3-8B and Llama-3-70B AI@Meta (2024)  [^5] https://huggingface.co/meta-llama/Meta-Llama-3-70B on 4×40G A100 GPUs loaded in fp16. We have tried to address the computation cost problem by introducing many effective acceleration or quantization packages like vllm  [^6] https://github.com/vllm-project/vllm, bitsandbytes  [^7] https://github.com/bitsandbytes-foundation/bitsandbytes, etc that are widely used to drive test time scaling law (Snell et al., 2024). As presented in Table 5, the results demonstrate the efficiency of our proposed UAlign method even though scaling on larger models.
Model Sampling Number Time Cost
Llama-3-8B 1 18 min
4 22 min
7 24min
10 25min
Llama-3-70B 1 1h 18min
4 1h 33min
7 1h 40min
10 2h 12min
Table 5: Time cost in different sampling numbers of UAlign on Llama-3-8B and Llama-3-70B.
In addition, the relatively low computation costs when sampling can be attributed that experiments are conducted on knowledge QA datasets. The answer spans are entity-level and each answer only needs to generate a few tokens. Since the output form is relatively simple, sampling ten times is sufficient and cost-saving to accurately fit the knowledge boundaries as presented in Sec. 4.4.
Furthermore, Test Time Scaling Law (Snell et al., 2024) has attracted much attention recently which proposed to consider allocating more computation resources in inference to generate high-quality responses. These LLMs’ self-generated data can be further used for LLM training to self-improve LLMs (Gulcehre et al., 2023). Many works validate that incorporating data multiply sampled on LLMs in inference can benefit LLMs for further improvements, which is a new trend for LLM training. In this way, our proposed UAlign provides a novel insight of the test time scaling law to represent knowledge boundaries by calculating uncertainty estimations on the sampled responses, and further explicitly leverages the uncertainty estimations for factuality alignment, heralding a promising view of test time scaling law. Although few additional computation costs are required, our proposed UAlign is still efficient to be utilized practically with significant reliability improvements.
B.3 Computation Cost of Inference of UAlign
Following B.2, we subsequently analyze the computation cost during inference of UAlign. Our proposed UAlign barely increases additional inference memory and time budget as follows.
First, uncertainty models also share the base LLMs with their respective plug-in LoRA modules with rank $r$=4. Additional parameters introduced only account for less than 1% of the base model parameters.
Second, uncertainty models only predict two tokens of uncertainty estimations in inference. We report the inference time cost on four test sets of vanilla Llama-3 which only generates the answer to the question and UAlign trained Llama-3 which predicts uncertainty estimations and then generates the answers on a single A100 GPU Card.
Dataset Time Cost
Vanilla ICL UAlign
TVQA 58 min 1h 6min
NQ-Open 28 min 32m in
SciQ 6 min 7 min
LSQA 5 min 6 min
Table 6: Inference time cost on four test sets of Llama-3 using vanilla ICL prompt-based and UAlign methods. Note the inference time costs on all the baseline methods in Sec. 3.3 are comparable to the vanilla ICL prompt-based baseline method.
Therefore, with the slight increase in additional memory and time cost in inference, UAlign significantly outperforms other baseline methods, demonstrating the reliability and efficiency on such tasks.
Appendix C Dataset Details